A compact fluorescent lamp (CFL) uses less energy than its equivalent incandescent bulb. How much will a person save if he/she uses a 20 W CFL instead of a 100 W incandescent lamp for 6 hours in 30 days? Electricity is priced at ₱ 8.17/kW.h.


1
Expert's answer
2022-04-19T10:00:16-0400

"m=0.001\cdot({100-20})\cdot 8.17\cdot 6\cdot 30=117.65."